Solve for the unknown interest rate in each of the following (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.):

  

Present Value Years Interest Rate Future Value
$ 270 4      %     $ 338
390 18               1,231
42,000 19               210,390
41,261 25               613,284
*** please need the right answer to this question ***Right ANSWER****
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1.

338=270(1+r/100)^4

(338/270)^(1/4)=(1+r/100)

(1+r/100)=1.0578

r=(1.0578-1)*100

=5.78%(Approx).

2.

1231=390(1+r/100)^18

(1231/390)^(1/18)=(1+r/100)

(1+r/100)=1.0659

r=(1.0659-1)*100

=6.59%(Approx).

3.

210390=42000(1+r/100)^19

(210390/42000)^(1/19)=(1+r/100)

(1+r/100)=1.0885

r=(1.0885-1)*100

=8.85%(Approx).

4.

613284=41261(1+r/100)^25

(613284/41261)^(1/25)=(1+r/100)

(1+r/100)=1.114

r=(1.114-1)*100

=11.40%(Approx).